OpenStudy (camerondoherty):

How do you make an equation from 2 points on a coordinate plane?

OpenStudy (anonymous):

You use the slope formula \(\huge\frac{y2-y1}{x2-x1}\) then you plug the slope and the points into the point-slope formula which is \(\huge(y-y1)=m(x-x1)\)

OpenStudy (camerondoherty):

oh ok Thank You